OpenStudy (anonymous):

Estimate three and three-fourteenths times five and eleven-twelfths by rounding each factor to the nearest whole number. fifteen eighteen twenty twenty-four

OpenStudy (shamil98):

\[\huge 3\frac{ 3 }{ 14 } \times 5\frac{ 11 }{ 12 }\] Rounding to the nearest whole number.. \[\huge 3 \times 6 = \] is your answer.